Orlando advertised for bids for the purchase of $3 million principal amount of Waste Water Revenue Bonds. Bonds will be delivered on April 1, 2021, and the interest will be paid on April 1st of the following years. The bonds mature as follows:
Maturity Date
Amount ($)
4/1/2025
100,000
4/1/2026
100,000
4/1/2027
100,000
4/1/2028
100,000
4/1/2029
200,000
4/1/2030
200,000
4/1/2031
250,000
4/1/2032
250,000
4/1/2033
250,000
4/1/2034
700,000
4/1/2035
750,000
The City received three competing bids for the Waste Water Revenue Bonds. The three offers are as follows:
From Rogue Securities:
• The City receives $3.5 million dollars
• The Interest Rates for the serial bonds with maturities:
o 2025 through 2030, 5.50 percent
o 2031 through 2035, 6.00 percent
From Johnson-Miller:
• The City Receives $3 million dollars
• The Interest Rates for the serial bonds with maturities:
o 2025 through 2027, 4.35 percent
o 2028 through 2032, 5.25 percent
o 2033 through 2035, 6.50 percent
From Shostak Corp:
• The City receives $2.9 million dollars
• The Interest rates for the serial bonds with maturities:
o 2025 to 2032, 5.75 percent
o 2033 to 2035, 6.25 percent
For each bid, compute the net interest cost (NIC) and the true interest cost (TIC). Which bid is more advantageous for the city?
Answer #1

Solution :

Note : the interest calculation here is pretty meticulous so for your understanding i have attached an excel formula sheet of all the 3 offers :

Net Interest Cost (NIC) : It is the notional interest cost that a bond issuer has to bear however it is not actual cost for issuer as this cost ignores time value of money which is taken care of in True interest cost(TIC).

Components of NIC calculation includes any discount, premium paid ( however this does not include time value of money)

***Formula NIC = (Total Interest Payments + Discount - Premium) / Number of Bond-Year Dollars***

True Interest Cost : It is similar to NIC but is more realistic in cost calculation since it includes time value of money and most widely used by bond issuers as it gives actual cost of interest

***Formula TIC = NIC / (1+ Discount rate)

  1. Interest Paid calculation of Rogue Securities

A 1 Bond value 3000000 2 Rogue Sec. 3 No. of bond year dollars Maturity Interest Paid 4 Year Interest - $C$1*$D$9 0 5 44287 6

NIC of Rogue Securities i.e marked green = 21.5175%

TIC ( since in question discount rate not given we assume it at 10% ), here no. of years = 15 i.e (2021-2035)
Thus, True interest cost = NIC / (1+0.1)^15 = 21.5175 / 0.239393 = 89.89 %


2.Interest Paid calculation of Jonson-miller :

G 1 Bond value 3000000 2 Jonson-miller 3 f bond vear dollars Maturity Year No Interest Rate Interest Paid 4 $C$1*SF$9 =sC$1*$

NIC of Jonson-miller i.e marked green = 17.0184%

TIC ( since in question discount rate not given we assume it at 10% ), here no. of years = 15 i.e (2021-2035)
Thus, True interest cost = NIC / (1+0.1)^15 = 17.0184 / 0.239393 = 71.09 %

3.Interest Paid calculation of Shostak-Corp :

2 Bond value 3000000 Shostak-Corp 4 Maturity Year No. of bond year dollars Interest Interest Paid sc$2*SH$10 6 44287 - -SC$2

NIC of Shostak-Corp i.e marked green = 22.4956%

TIC ( since in question discount rate not given we assume it at 10% ), here no. of years = 15 i.e (2021-2035)
Thus, True interest cost = NIC / (1+0.1)^15 = 22.4956 / 0.239393 = 93.96 %

Conclusion : Offer 1 is the best as it gives a high return to bond holders of 21.5% along with highest maturity value of $3.5 million